# TinkerXR: In-Situ, Reality-Aware CAD and 3D Printing Interface for Novices

### v3

@@ 4.3.2. Drag and Drop and Adjustments on Build Plane
Users can pick up the models they created, and drag them to the virtual printer twin with the movement handle, where the printer will be highlighted (Figure 12a). After a model is dropped over the printer twin, the duplicate of the objects comprised by the model will be placed inside the printer (Figure 12b). The original objects will be moved to their initial position. After the object is placed inside the printer, users can adjust the size, orientation, and position of the object (Figure 12c). Note that a print can not be initiated when the objects go outside the boundaries of the build area.
4.3.3. Preparing for Printing and Slicing
@@
The sliced model is visualized in the virtual printer (Figure 13b and 13c), where the main walls are displayed in gray and support structures in cyan (tree support selected). The visualization allows users to review the slicing results and identify areas requiring adjustment before starting the print. If the selected model exceeds the printer’s build area, the system notifies users by displaying a ”too large for printer” indicator above the virtual printer twin (Figure 13d). This ensures that users can modify their design or scale it appropriately before initiating the print.
@@
After the printing request is sent to the virtual printer twin object, if no server IP is given, the STL model of the object in the virtual printer will be saved locally. If the server IP is given while the printer IP is not, this STL model and sliced G-code file will be saved on the remote server. However, if both IPs are given, the printing process will start after slicing. The headset monitors those steps. The server IP is the same as the computer the server program is running on, and the printer IP can be obtained from the printer after it has been started and connected to the internet. Note that the printer should be accessible from the server computer if they do not reside on the same subnet.
@@
The server manages multiple slicers to serve the XR headset clients. An example slicer configuration for testing has been created for Ultimaker S3121212https://ultimaker.com/3d-printers/s-series/ultimaker-s3/ printer with options: Enable Support, Support Structure, Layer Height, Infill Density, and Infill Pattern (Figure 13a). More slicers can be implemented by adding them to the slicer_options.json file. The slicer options are completely modifiable and could be increased with ease as the run command is also read from the JSON. How the slicer should work is completely in the hands of the programmer as the command could be easily manipulated by the flexible replacement tags that have been noted as options. The only required arguments that a slicer should take are output and input file arguments, which use gcode_path and stl_path, respectively.
@@
A client can make multiple requests on the server. When a /slice GET request is sent to the server, it returns the available slicers and their options. On the other hand, when a /slice POST request is sent to the server, the STL file is loaded on the appropriate slicer, and the G-code file of the target model is obtained. The server parses the options selected by the user and passes them to the slicer. After the slicing operation has been completed, the resulting G-code file is visualized in the virtual printer. This draft object is shown in different colors for the main and the support parts. After that, the printing process can start. Note that one server can handle multiple printers, and one headset can connect to various servers. The only necessary action is to enter IP addresses.
4.3.4. Printing and Monitoring
@@
After the /print POST request is sent to the virtual printer twin object, the server will connect and get access permission to the 3D printer. This permission is only necessary the first time as the server will save the token for accessing the printer, and access permission will not be necessary until the token has expired. If access to the 3D printer is granted, the server will send a printing request with the G-code file to the printer. The headset can obtain the print process information by sending the /print GET request. The user can pause or abort the printing process by sending a PUT request to the /print URL if necessary. This request supports three different commands: continue, pause, and stop, which can be sent via the buttons next to the printer twin (Figure 12d).
@@ 5. Applications
Figure 14. Designs created via TinkerXR on the left; their fabricated results on the right, placed in their intended space. These are (a) phone holder, (b) coaster, (c) pen holder, (d) candy bowl, (e) finger skateboard obstacle, (f) figurine stand, (g) key hanger, (h) door wedge, (i) plant pot hanger, (j) lamp shade.
[…]
Ten images labeled (a), (b), (c), (d), (e), (f), (g), (h), (i), and (j) show ten different 3D printed results next to their virtual rendered versions. The image (a) has a phone holder with a smartphone on top. In the image (b), there is a drink coaster with a cup on top of it. The image (c) has a pen holder with colorful markers inside it. In the image (d), there is a candy bowl with packaged candy inside. In the image (e), there is a ramp and a grinding rail with a finger skateboard next to it. In image (f), a figurine stand is mounted to a wall with a small figurine. In the image (g), there is a wall key hanger with two hooks and a bunch of keys hanging down those hooks. In the image (h), there is a door wedge holding a door from closing. In image (i), there is a plant pot carrying a ring connected to ropes that are hanging from the ceiling. In the image (j), there is a spherical light bulb shade covering a light bulb mounted to the ceiling.
@@ 5.2. Wall Workspace
For the wall workspace, two objects were designed: a figurine stand and a key hanger (Figure 14f and 14g). The figurine stand allows users to display small decorative items without adding clutter, created using the ruler tool for a fitting hole. The key hanger offers practical functionality with aligned hooks using snap-to-grid. The models are snapped to the wall while designing for a better sense of depth, helping users create visually integrated solutions.
@@ 5.3. Other Workspaces
Beyond the more common desk and wall spaces, TinkerXR can also be applied to other environments, e.g., ground, ceiling, doors, windows, screens, couches, and more. On the ground, a door wedge was designed to hold doors securely in place (Figure 14h), benefiting from real-time scaling for accurate size.
For the ceiling, we created two distinct designs: a plant pot hanger and a light shade (Figure 14i and 14j). The plant pot hanger demonstrates the system’s ability to accommodate objects that require support and suspension, such as using ”ropes” for hanging, a practical solution for overhead applications. The light shade, on the other hand, highlights the system’s capability to handle complex and abstract 3D modeling. Its artistic design illustrates the system’s flexibility in creating intricate structures on a headset without technical limitations. These examples show the adaptability to less conventional design spaces and orientations.
@@ 6.1. Methods
Participants. We recruited ten participants (ages 23-29; 9 male, 1 female). Four participants had little experience with CAD software, and six had no previous CAD experience. Additionally, three participants had little experience with AR systems, and seven had no experience with AR systems. Finally, four have used 3D printers in some parts of their life, while six have never used a 3D printer.
@@
Figure 15. Designs created by the participants in the user study. Objects are colored based on their task number (a: wall-mounted key hanger, b: cable management, c: toiletries holder) and the system used for designing the models.
[…]
@@
Task & Procedure. For the evaluation, we conducted a within-subjects comparative study to assess TinkerXR against Tinkercad. Each participant completed modeling tasks using both tools, with the order of tool usage counterbalanced across participants to mitigate order effects (e.g., half of the participants used TinkerXR first, followed by Tinkercad, while the other half did the opposite). We showed both systems briefly, and participants completed a 3-minute warm-up before starting the tasks to familiarize themselves with the tools. Participants were assigned three modeling tasks, each tailored to a specific context and accompanied by explicit design requirements: #1 Wall-mounted key hanger: Participants created a key hanger with two hooks narrow enough to accommodate keychain rings. This task emphasized designing small, precise features and required the use of a vertical wall as the design context. #2 Cable management: Participants designed a holder to organize cables on a desk. The design had to securely fit the provided cables, ensuring they would not fall, and the organizer needed to remain stable when attached to the desk. #3 Toiletries holder: Participants designed a holder for a toothpaste tube and toothbrush, intended for use in a bathroom. They were provided with physical objects representing the toiletries and used the reference object functionality to ensure the holder dimensions aligned with the objects. For both conditions, participants were allowed to use a physical ruler to measure physical dimensions, although they were also allowed to use the digital AR ruler in the TinkerXR condition.
@@
Participants engaged with TinkerXR using a Quest 3 headset and hand interactions. For TinkerXR, they were instructed to prepare and start their prints on a virtual twin of the 3D printer; while for the Tinkercad condition, they used Cura [^13] https://ultimaker.com/software/ultimaker-cura/’s desktop UI for 3D printing preparation. To save time, participants did not wait for the physical prints to finish before moving on to the next task. The printer slicing sequence was only carried out for the first task due to the same reason. Each modeling session was followed by a survey to gather qualitative and quantitative feedback on the tools, focusing on usability, efficiency, and overall satisfaction.
@@
Measures. We used both quantitative and qualitative metrics to evaluate the system. We recorded the time required to complete the individual 3D design tasks as a measure of overall system performance. For Task 1, we also measured the time required to transfer the finished design into the 3D printing environment and prepare it for fabrication, i.e., slicing. We measured this duration until the participant hit the print button. After completing the tasks, participants were asked to complete standardized questionnaires, including the System Usability Scale (SUS) (Brooke, 1996), which we extended with additional questions related to the immersive experience, and the NASA Task Load Index (NASA-TLX) (Hart and Staveland, 1988). We collected qualitative feedback through open-ended questions regarding their overall experience.
@@ 6.2.1. Observations and Analysis of the Fabricated Artifacts
Figure 15 shows the three types of objects designed and fabricated by the participants. Each design task had specific requirements to ensure functionality and usability, and all participants successfully met these requirements. For the key hanger, the design required at least two hooks to hold keychains. The hooks had to be narrow enough to accommodate standard keychain rings. For the cable management, participants needed to design a holder that securely fit at least two cables of provided diameters, ensuring the cables would not slip out. Additionally, the design had to accommodate the thickness of the desk to remain stable when attached. For the toiletries holder, the design required at least two compartments or holes to hold the provided toothbrush and toothpaste tube. These compartments had to match the given dimensions to ensure a stable hold. All models were validated by checking if they adhered to dimensional requirements and functionality constraints defined for each task.
A notable insight was that cube primitive as a solid and cylinder geometry as a hole were mostly used objects for both the TinkerXR and Tinkercad conditions. During TinkerXR usage, snap-to-grid was heavily utilized. All of the participants used the physical ruler for the wall-mounted key hanger and cable management tasks while designing with Tinkercad, while all of them used the reference object functionality of TinkerXR for the toiletries holder while designing with TinkerXR. Participants relied heavily on depth occlusion and the highlighted projection of the objects on the grid to understand the location and depth of the objects. Seen in Figure 15, the outcomes that have taken the most time on average of two systems were the key hanger by P5 (18m 45s), the toiletiers holder by P5 (15m 31s), and the key hanger by P9 (14m 35s).
All of the resulting objects with TinkerXR vs. Tinkercad were comparable, i.e., none of the users created designs that differed deeply depending on the tool used. This verifies that TinkerXR was able to offer similar 3D CAD capabilities despite the newly introduced immersive AR environment and not being an established commercial solution.
@@ 6.2.2. Time
Across all three design tasks, we found no statistically significant difference in task completion time using TinkerXR compared to Tinkercad: Task 1 (t = 0.272, p = 0.792), Task 2 (t = 0.272, p = 0.792), Task 3 (t = -1.769, p = 0.110). This suggests that TinkerXR provides an immersive design experience without a substantial time cost. However, future work may explore this tradeoff in larger or more varied user populations, e.g., for experts as well. Fig. 16 shows the results for all tasks. When it comes to the slicing process for Task 1, the 3D printing preparation took less time when participants utilized the TinkerXR system. A paired t-test confirmed significantly shorter completion times in the TinkerXR condition for Task 1 (t = 3.747, p = 0.004).
Figure 16. Recorded task completion times (in seconds) for participants using Tinkercad and TinkerXR across three design tasks and one slicing task. No statistically significant difference in completion times was found between the two systems for the design tasks. For the slicing task, participants completed the process significantly faster with TinkerXR compared to Tinkercad (t = 3.747, p = 0.004).
[…]
Result of the recorded times of the participants.
@@ 6.2.3. Surveys
For the standard SUS survey, we obtained a score of 90%, which shows that TinkerXR system is within the A grade (90% or above), while Tinkercad achieved a score of 85%, which places it within the B grade (from 80% to 90%)  (Sauro and Lewis, 2016). As shown in Fig. 17, the TinkerXR system has superior scores across statements such as Q3, Q9, and Q10. These findings support that TinkerXR was easier to use for users in terms of coherency. The additional questions related to immersive experience show that TinkerXR helped users to have a more natural and hands-on designing procedure.
Figure 17. Results of the extended SUS survey comparing Tinkercad and TinkerXR. User ratings are shown on a Likert scale from 1 (strongly disagree) to 5 (strongly agree). TinkerXR demonstrates generally higher ratings for immersive design features.
[…]
Results of the SUS survey.
The analysis of the NASA-TLX survey with a paired t-test reveals significant differences between Tinkercad and TinkerXR across multiple workload dimensions. Participants reported lower mental demand when using TinkerXR compared to Tinkercad (t = 3.857, p = 0.004), indicating that the immersive AR environment facilitated easier comprehension and interaction during design tasks. In contrast, TinkerXR exhibited higher physical demand (t = -5.658, p <0.001) and effort (t = -2.684, p = 0.025), likely due to the reliance on hand gestures and the physical interaction required with the AR system compared to the mouse-and-keyboard interface of Tinkercad. However, temporal demand, performance, and frustration levels were comparable between the two tools, indicating that the usability benefits of TinkerXR may mitigate the added physical strain.
@@
Figure 18. Results of the NASA-TLX survey comparing Tinkercad and TinkerXR across six workload dimensions. TinkerXR shows lower mental demands and higher perceived performance, while Tinkercad exhibits lower physical demand. Performance, temporal, and frustration scores are comparable between the two tools.
[…]
Result of the NASA-TLX survey.

### v4

@@ 4.3.2. Drag and Drop and Adjustments on Build Plane
Users can pick up the models they created, and drag them to the virtual 3D printer twin with the movement handle, where the printer will be highlighted (Figure 9a). After a model is dropped over the printer twin, the duplicate of the objects comprised by the model will be placed inside the printer twin (Figure 9b). The original objects will be moved to their initial position. After the object is placed inside the printer twin, users can adjust the size, orientation, and position of the object (Figure 9c). Note that a print can not be initiated when the objects go outside the boundaries of the build area.
4.3.3. Preparing for 3D Printing and Slicing
@@
The sliced model is visualized in the virtual 3D printer (Figure 10b and 10c), where the main walls are displayed in gray and support structures in cyan (tree support selected). The visualization allows users to review the slicing results and identify areas requiring adjustment before starting the print. If the selected model exceeds the 3D printer’s build area, the system notifies users by displaying a ”too large for printer” indicator above the virtual printer twin (Figure 10d). This ensures that users can modify their design or scale it appropriately before initiating the print.
@@
After the printing request is sent through the buttons next to the virtual 3D printer twin object, if no server IP is given, the STL model of the object in the virtual printer will be saved locally. If the server IP is given while the printer IP is not, this STL model and sliced G-code file will be saved on the remote server. However, if both IPs are given, the printing process will start after slicing. The headset monitors those steps. The server IP is the same as the computer the server program is running on, and the printer IP can be obtained from the printer after it has been started and connected to the internet. Note that the printer should be accessible from the server computer if they do not reside on the same subnet.
@@
The server manages multiple slicers to serve the XR headset clients. An example slicer configuration for testing has been created for Ultimaker S3131313https://ultimaker.com/3d-printers/s-series/ultimaker-s3/ printer with options: Enable Support, Support Structure, Layer Height, Infill Density, and Infill Pattern (Figure 10a). More slicers can be implemented by adding them to the slicer_options.json file. The slicer options are completely modifiable and could be increased with ease as the run command is also read from the JSON. How the slicer should work is completely in the hands of the programmer as the command could be easily manipulated by the flexible replacement tags that have been noted as options. The only required arguments that a slicer should take are output and input file arguments, which use gcode_path and stl_path, respectively.
@@
A client can make multiple requests on the server. When a /slice GET request is sent to the server, it returns the available slicers and their options. On the other hand, when a /slice POST request is sent to the server, the STL file is loaded on the appropriate slicer, and the G-code file of the target model is obtained. The server parses the options selected by the user and passes them to the slicer. After the slicing operation has been completed, the resulting G-code file is visualized in the virtual 3D printer. This draft object is shown in different colors for the main and the support parts. After that, the 3D printing process can start. Note that one server can handle multiple printers, and one headset can connect to various servers. The only necessary action is to enter IP addresses.
4.3.4. 3D Printing and Monitoring
@@
After the /print POST request is sent to the virtual printer twin object, the server will connect and get access permission to the 3D printer. This permission is only necessary the first time as the server will save the token for accessing the printer, and access permission will not be necessary until the token has expired. If access to the 3D printer is granted, the server will send a printing request with the G-code file to the printer. The headset can obtain the print process information by sending the /print GET request. The user can pause or abort the printing process by sending a PUT request to the /print URL if necessary. This request supports three different commands: continue, pause, and stop, which can be sent via the buttons next to the 3D printer twin (Figure 9d).
@@ 5. Applications
Figure 11. Designs created via TinkerXR on the left; their fabricated results on the right, placed in their intended space. These are (a) phone holder, (b) coaster, (c) pen holder, (d) candy bowl, (e) finger skateboard obstacle, (f) figurine stand, (g) key hanger, (h) door wedge, (i) plant pot hanger, (j) lamp shade.:
@@ 5.2. Wall Workspace
For the wall workspace, two objects were designed: a figurine stand and a key hanger (Figure 11f and 11g). The figurine stand allows users to display small decorative items without adding clutter, created using the ruler tool for a fitting hole. The key hanger offers practical functionality with aligned hooks using snap-to-grid. The models are snapped to the wall while designing for a better sense of depth, helping users create visually integrated solutions.
@@ 5.3. Other Workspaces
Beyond the more common desk and wall spaces, TinkerXR can also be applied to other environments, e.g., ground, ceiling, doors, windows, screens, couches, and more. On the ground, a door wedge was designed to hold doors securely in place (Figure 11h), benefiting from real-time scaling for accurate size. It was also crucial to have a feature where users can work around and build with existing objects in mind, so the coordinate system does not only depend on flat surfaces to stand on, but also considers the real-world objects they will interact with. The reference objects and the ability to modify the coordinate system based on any vertex of a reference object are directed towards this solution.
For the ceiling, we created two distinct designs: a plant pot hanger and a light shade (Figure 11i and 11j). The plant pot hanger demonstrates the system’s ability to accommodate objects that require support and suspension, such as using ”ropes” for hanging, a practical solution for overhead applications. The light shade, on the other hand, highlights the system’s capability to handle complex and abstract 3D modeling. Its artistic design illustrates the system’s flexibility in creating intricate structures on a headset without technical limitations. These examples show the adaptability to less conventional design spaces and orientations.
@@ 6.1. Methods
We recruited ten participants (ages 23-29; 9 male, 1 female). Four participants had little experience with CAD software, and six had no previous CAD experience. Additionally, three participants had little experience with AR systems, and seven had no experience with AR systems. Finally, four have used 3D printers in some parts of their life, while six have never used a 3D printer.
@@
Figure 12. Designs created by the participants in the user study. Objects are colored based on their task number (a: wall-mounted key hanger, b: cable management, c: toiletries holder) and the system used for designing the models.:
Task & Procedure
@@
For the evaluation, we conducted a within-subjects comparative study to assess TinkerXR against Tinkercad. Each participant completed modeling tasks using both tools, with the order of tool usage counterbalanced across participants to mitigate order effects (e.g., half of the participants used TinkerXR first, followed by Tinkercad, while the other half did the opposite). We showed both systems briefly, and participants completed a 3-minute warm-up before starting the tasks to familiarize themselves with the tools. Participants were assigned three modeling tasks, each tailored to a specific context and accompanied by explicit design requirements: #1 Wall-mounted key hanger: Participants created a key hanger with two hooks narrow enough to accommodate keychain rings. This task emphasized designing small, precise features and required the use of a vertical wall as the design context. #2 Cable management: Participants designed a holder to organize cables on a desk. The design had to securely fit the provided cables, ensuring they would not fall, and the organizer needed to remain stable when attached to the desk. #3 Toiletries holder: Participants designed a holder for a toothpaste tube and toothbrush, intended for use in a bathroom. They were provided with physical objects representing the toiletries and used the reference object functionality to ensure the holder dimensions aligned with the objects. For both conditions, participants were allowed to use a physical ruler to measure physical dimensions, although they were also allowed to use the digital AR ruler in the TinkerXR condition.
@@
Participants engaged with TinkerXR using a Quest 3 headset and hand interactions. For TinkerXR, they were instructed to prepare and start their prints on a virtual twin of the 3D printer; while for the Tinkercad condition, they used Cura [^14] https://ultimaker.com/software/ultimaker-cura/’s desktop UI for 3D printing preparation. To save time, participants did not wait for the physical prints to finish before moving on to the next task. The 3D printer slicing sequence was only carried out for the first task due to the same reason. Each modeling session was followed by a survey to gather qualitative and quantitative feedback on the tools, focusing on usability, efficiency, and overall satisfaction.
Measures
@@
We used both quantitative and qualitative metrics to evaluate the system. We recorded the time required to complete the individual 3D design tasks as a measure of overall system performance. For Task 1, we also measured the time required to transfer the finished design into the 3D printing environment and prepare it for fabrication, i.e., slicing. We measured this duration until the participant hit the print button. After completing the tasks, participants were asked to complete standardized questionnaires, including the System Usability Scale (SUS) (Brooke, 1996), which we extended with additional questions related to the immersive experience, and the NASA Task Load Index (NASA-TLX) (Hart and Staveland, 1988). We collected qualitative feedback through open-ended questions regarding their overall experience.
@@ 6.2.1. Observations and Analysis of the Fabricated Artifacts
Figure 12 shows the three types of objects designed and fabricated by the participants. Each design task had specific requirements to ensure functionality and usability, and all participants successfully met these requirements. For the key hanger, the design required at least two hooks to hold the keychains. The hooks had to be narrow enough to accommodate standard keychain rings. For cable management, participants needed to design a holder that would securely fit at least two cables of the diameters provided, to ensure that the cables would not slip out. The design also had to accommodate the thickness of the desk to remain stable when attached. For the toiletries holder, the design required at least two compartments or holes to hold the provided toothbrush and toothpaste tube. These compartments had to match the given dimensions to ensure a stable hold. All models were validated by checking if they adhered to the defined dimensional requirements and functionality constraints.
A notable insight was that cube primitive as a solid and cylinder geometry as a hole were mostly used objects for both the TinkerXR and Tinkercad conditions. During TinkerXR usage, snap-to-grid was heavily utilized. All participants used the physical ruler for the wall-mounted key hanger and cable management tasks while designing with Tinkercad, while all used the reference object functionality for the toiletries holder while designing with TinkerXR. Participants relied heavily on depth occlusion and the highlighted projection of the objects on the grid to understand the location and depth of the objects. Seen in Figure 12, the outcomes that have taken the most time on average of the two systems were the key hanger by P5 (18m 45s), the toiletiers holder by P5 (15m 31s), and the key hanger by P9 (14m 35s). We note that, as seen in Figure 12, the majority of the resulting objects created with TinkerXR vs. Tinkercad were comparable, i.e., none of the users created designs that differed deeply depending on the tool used. This verifies that TinkerXR was able to offer similar 3D CAD capabilities despite the newly introduced immersive AR environment and not being an established commercial solution as a research artifact.
Figure 13. Results of the extended SUS survey comparing Tinkercad and TinkerXR. User ratings are shown on a Likert scale from 1 (strongly disagree) to 5 (strongly agree). TinkerXR demonstrates generally higher ratings for immersive design features.:
@@ 6.2.2. Time
Across all three design tasks, we found no statistically significant differences in task completion time using TinkerXR compared to Tinkercad: Task 1 (t = 0.272, p = 0.792), Task 2 (t = 0.272, p = 0.792), Task 3 (t = -1.769, p = 0.110). This suggests that TinkerXR provides an immersive design experience without a substantial time cost. However, future work may explore this tradeoff in larger or more varied user populations, e.g., for experts as well. When it comes to the slicing process for Task 1, the 3D printing preparation took less time when participants utilized TinkerXR. A paired t-test confirmed significantly shorter completion times in the TinkerXR condition for Task 1 (t = 3.747, p = 0.004).
@@ 6.2.3. Surveys
For the standard SUS survey, we obtained a score of 90%, which shows that TinkerXR system is within the A grade (90% or above), while Tinkercad achieved a score of 85%, which places it within the B grade (from 80% to 90%)  (Sauro and Lewis, 2016). As shown in Fig. 13, TinkerXR has higher scores across statements such as Q3, Q9, and Q10. These findings support that TinkerXR was easier to use for users in terms of coherence. The additional questions related to immersive experience show that TinkerXR helped experience a more natural and hands-on design procedure.
Analysis of the NASA-TLX survey with a paired t-test reveals significant differences between Tinkercad and TinkerXR across multiple workload dimensions. Participants reported lower mental demand when using TinkerXR compared to Tinkercad (t = 3.857, p = 0.004), indicating that the immersive AR environment facilitated easier comprehension and interaction during design tasks. In contrast, TinkerXR exhibited higher physical demand (t = -5.658, p <0.001) and effort (t = -2.684, p = 0.025), likely due to the dependence on hand gestures and the physical interaction required with the AR system compared to the mouse-and-keyboard interface of Tinkercad. However, temporal demand, performance, and frustration levels were comparable between the two, indicating that the usability benefits of TinkerXR may mitigate the added physical strain.
@@
Figure 14. Results of the NASA-TLX survey comparing Tinkercad and TinkerXR across six workload dimensions. TinkerXR shows lower mental demands and higher perceived performance, while Tinkercad exhibits lower physical demand. Performance, temporal, and frustration scores are comparable between the two tools. :